A box contain 74 brass washers ,86 steel washers and 40 aluminium washers three washers are down at random from the box without replacement .Determine the probability that all three are steel washers

Answer by Fombitz(32388)

74 B, 86 S, 40 A gives 200 total washers.
P(1st S washer)=86/200
P(2nd S washer)=85/199
P(3rd S washer)=84/198
P(3 S washer)=(86/200)(85/199)(84/198)=5117/65670